OpenStudy (barrelracing):

The graph of f(x) = 2x + 1 is shown below. Explain how to find the average rate of change between x = 0 and x = 3.

OpenStudy (caozeyuan):

rate of chane is slope, since slope is 2, that's your answer

OpenStudy (photon336):

evaluate f(x) at x = 0 and x = 3 and then maybe use the slope formula. \[\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} } = m_{average}\]

OpenStudy (caozeyuan):

rate of change is dy/dx, this is linear, so it is just the slope

OpenStudy (barrelracing):

so I'm confused how do I find the average rate of change

OpenStudy (caozeyuan):

look at the slope

OpenStudy (caozeyuan):

rate of change gets complicated beyond linear, but this is linear so no worry

OpenStudy (photon336):

first find what the y values are. \[f(0) = 2(0)+1 = 1 \] \[f(3) = 2(3)+1 = 7 \] 0,1 and 3,7 since it's just a line you can use the slope formula \[m = \frac{ y_{2}-y_{1} }{ x_{2}-x_{1}}\] then use the formula above to find the slope between those two points and that's the average rate of change. you could use derivatives too.

OpenStudy (photon336):

\[y_{2} = 7, y_{1} = 1, x_{2} = 3, x_{1} = 0 \] \[\frac{ 7-1 }{ 3-0 } = \frac{ 6 }{ 3 } \]

OpenStudy (photon336):

since it's a line you could avoid all this and just do what @caozeyuan said but that's if you haven't learned derivatives yet but that's just a technical term for slope.

OpenStudy (barrelracing):

thank u both.

OpenStudy (photon336):

@Barrelracing if you wanted you could take the derivative

OpenStudy (photon336):

\[y = m(x) + b\] \[\frac{ dY }{ dx } = m \]

OpenStudy (barrelracing):

well didn't u just provide the answer?

OpenStudy (photon336):

in case you wanted a faster way to do it

OpenStudy (photon336):

\[\frac{ dY }{ dX }, 2x+1 = 2x ^{1-1} = 2x ^{0} = 2\]